OpenStudy (anonymous):

I will meadel and fan!!!!!!!!!!!!!!!!! round the mixednumber to the nearest 1/2. 100 1/3 A. 100 B. 100 1/2 C. 101

OpenStudy (anonymous):

100 1/3=300+1/100=301/3=100.3334

OpenStudy (anonymous):

THATS NOT HELPFULALL I SEE IS BLLAA BLAA BLAA I GREEN KNOWS HOW TO BREAK IT DOWN

OpenStudy (anonymous):

UNLESS U KNOW HOW TO BREAK IT DOWN IAM NOT LISTENING

OpenStudy (igreen):

100 1/3 is approximately 100.33 Is 100.33 closer to 100.5 or 100?

OpenStudy (anonymous):

100

OpenStudy (igreen):

\(100.33 - 100 = 0.33\) \(100.5 - 100.33 = 0.17\) Which is bigger, \(0.17\) or \(0.33\)?

OpenStudy (anonymous):

0.33

OpenStudy (igreen):

Right, so that means 100 is wrong, so what's our answer going to be?

OpenStudy (anonymous):

100 1/2

OpenStudy (anonymous):

HELP!!!!!!!!!!

OpenStudy (anonymous):

is it right!!!!!!!!!!!!!!

OpenStudy (anonymous):

are u there

OpenStudy (igreen):

Yes, 100 1/2 is correct.

OpenStudy (anonymous):

thanks